# Fine Tuning vs. Retrieval Augmented Generation for Less Popular Knowledge

### v3

@@ 5. Results
Comparison of fine tuning methods. To study the effect of fine-tuning method, we investigate four LMs: three from the FlanT5 family (encoder-decoder models) and StableLM2 (a decoder-only model). These models are chosen as their parameters are less than 2B, allowing us to perform full fine-tuning with resources we had at our disposal. Table 2 shows that PEFT leads to smaller gains in the +FT-RAG setup compared to full FT in most cases, yet it significantly improves accuracy in the +FT+RAG setup. This suggests that PEFT enables the LM to maintain its reasoning abilities based on the provided prompts. Based on this observation, we selected PEFT as the fine-tuning method for subsequent experiments.
Comparison of data augmentation methods. The E2E method generates over 12 times more QAs than the prompting method, while the prompt-based method generates higher quality QAs. The results in Table 2 show that fine-tuned models trained on prompt-generated data outperform those trained on E2E-generated data. This highlights the significance of synthetic data quality over quantity. As a result, for subsequent experiments, we use QAs generated by the prompt method.
Comparison of FT and RAG with LMs of different type and size. Choosing PEFT as the FT method and the Prompt method for QA generation, we extended our experiments to 12 LMs using the ideal retriever for RAG. Table 3 presents the results. The findings indicate that while FT enhances the answer generator’s performance, it alone cannot match or approach the performance of RAG. However, combining FT with RAG yields the best results for smaller models (up to 3B parameters). Conversely, for larger models (from 7B to 11.3B parameters), combining FT with RAG degrades performance. This suggests that although FT injects knowledge into LMs (as seen when comparing -FT-RAG with +FT-RAG), it diminishes reasoning abilities of larger LMs. Interestingly, since smaller LMs’ accuracy improves with FT, the best result for +FT+RAG is achieved by StableLM2, which has only 1.6B parameters. Another observation is that decoder-only models perform better than encoder-decoder (Flan-T5) models of similar size.
@@
Effect of retrieval models on RAG. To assess the effect of retrieval models, we first calculate the overall Recall score for all retriever models, considering the first paragraph of each Wikipedia page as the answer-containing document. The results are shown in Table 4, demonstrating that DPR and BM25+DPR models significantly outperform BM25 and Contriever across all datasets, while DPR and BM25+DPR perform on par with other. Table 5 presents the QA accuracy of RAG with different retrieval models. Following Mallen et al. (2023), we use the top-ranked document for the RAG prompt. The results indicate a direct correlation between the performance of the retrieval model and the overall QA accuracy, underscoring the significant impact of the retrieval model on the effectiveness of the downstream task for both vanilla and fine-tuned LMs. As performance of RAG with DPR and BM25+DBPR is comparable, we use DPR as the retrieval model for subsequent experiments.

### v5

@@ 5. Results
Comparison of fine tuning methods. To study the effect of fine-tuning method, we investigate four LMs: three from the FlanT5 family (encoder-decoder models) and StableLM2 (a decoder-only model). These models are chosen as their parameters are less than 2B, allowing us to perform full fine-tuning with the resources we had at our disposal. Table 2 shows that PEFT leads to smaller gains in the +FT-RAG setup compared to full FT in most cases, yet it significantly improves accuracy in the +FT+RAG setup. This suggests that PEFT enables the LM to maintain its reasoning abilities based on the provided prompts. Based on this observation, we selected PEFT as the fine-tuning method for subsequent experiments.
Comparison of data augmentation methods. The E2E method generates over 12 times more QAs than the prompting method, while the prompt-based method generates higher-quality QAs. The results in Table 2 show that fine-tuned models trained on prompt-generated data outperform those trained on E2E-generated data. This highlights the significance of synthetic data quality over quantity. As a result, for subsequent experiments, we use QAs generated by the prompt method.
Comparison of FT and RAG with LMs of different type and size. Choosing PEFT as the FT method and the Prompt method for QA generation, we extended our experiments to 12 LMs using the ideal retriever for RAG. Table 3 presents the results. The findings indicate that while FT enhances the answer generator’s performance, it alone cannot match or approach the performance of RAG. However, combining FT with RAG yields the best results for smaller models (up to 3B parameters). Conversely, for larger models (from 7B to 11.3B parameters), combining FT with RAG degrades performance. This suggests that although FT injects knowledge into LMs (as seen when comparing -FT-RAG with +FT-RAG), it diminishes the reasoning abilities of larger LMs. Interestingly, since smaller LMs’ accuracy improves with FT, the best result for +FT+RAG is achieved by StableLM2, which has only 1.6B parameters. Another observation is that decoder-only models perform better than encoder-decoder (Flan-T5) models of similar size.
@@
Effect of retrieval models on RAG. To assess the effect of retrieval models, we first calculate the overall Recall score for all retriever models, considering the first paragraph of each Wikipedia page as the answer-containing document. The results are shown in Table 4, demonstrating that DPR and BM25+DPR models significantly outperform BM25 and Contriever across all datasets, while DPR and BM25+DPR perform on par with each other. Table 5 presents the QA accuracy of RAG with different retrieval models. Following Mallen et al. (2023), we use the top-ranked document for the RAG prompt. The results indicate a direct correlation between the performance of the retrieval model and the overall QA accuracy, underscoring the significant impact of the retrieval model on the effectiveness of the downstream task for both vanilla and fine-tuned LMs. Since the performance of RAG with DPR and BM25+DPR is comparable, we use DPR as the retrieval model for subsequent experiments.